OpenStudy (anonymous):

How can I factor this trinomial -8x2 – 10x +3?

OpenStudy (anonymous):

use middle term factorisation

OpenStudy (anonymous):

Let the coefficient of x^2(i.e.8)=a coefficient of x=b n the constant term be c therefore a=8 b=10 c=3 find two numbers m and n such that m+n=b and m*n=a*c

OpenStudy (anonymous):

@Rishi123 is using the standard form of a linear equation Ax^2 + Bx + C = 0. That is where he is getting the a, b, and c.

OpenStudy (anonymous):

The polynomial can be factored by guess and check method or using the quadratic formula. Would you like me to walk you through it?